A car with good tires on a dry road can decelerate at about 5.5 m/s2 when braking. Suppose a car is initially traveling at 62 mi/h.

a) How much time does it take the car to stop?

b) what is the stopping distance?

Please show step by step!
Answer #1

here,

deaccelerates , a = - 5.5 m/s^2

initial speed of car , u = 62 mi/h

u = 27.72 m/s

a)

let the time taken be t

using first equation of motion

v = u + a * t

0 = 27.72 - 5.5 * t

t = 5.04 s

the time taken is 5.04 s

b)

the stopping distance , s = u^2 /(2a)

s = 27.72^2 /(2*5.5) m

s = 69.85 m